OpenStudy (anonymous):

solve algebraically. y=x^2 and 7x-y=12

OpenStudy (gw2011):

Substitute y=x^2 into 7x-y=12 7x-x^2=12 0=x^2-7x+12 This equation can be re-written as follows: x^2-7x+12=0 Factor this equation (x-4)(x-3)=0 x=4 and y=16 Because y=x^2=(4)^2=16 x=3, y=9 To check the answer substitute x=4 and y=16 into 7x-y=12 and substitute x=3 and y=9 into 7x-y=12 and in both cases you will get the equation 7x-y equal to 12